OpenStudy (itsmichelle29):

What is the sum of a 6-term geometric series if the first term is 21 and the last term is 1,240,029? 1,395,030 1,461,460 1,527,890 1,594,320

OpenStudy (campbell_st):

well you need to find the value of the common ratio. you can do this by using the formula \[a_{n} = a_{1} \times r^{n - 1}\] so using the information \[1240029 = 21\times r^5\] you need to solve this equation for r then the sum of 6 terms is \[s_{6} = \frac{21( 1 - r^6)}{1 - r}\]